# PROPHET: An Inferable Future Forecasting Benchmark with Causal Intervened Likelihood Estimation

### v1

@@ Abstract
Predicting future events stands as one of the ultimate aspirations of artificial intelligence. Recent advances in large language model (LLM)-based systems have shown remarkable potential in forecasting future events, thereby garnering significant interest in the research community. Currently, several benchmarks have been established to evaluate the forecasting capabilities by formalizing the event prediction as a retrieval-augmented generation (RAG)-and-reasoning task. In these benchmarks, each prediction question is answered with relevant retrieved news articles. However, because there is no consideration on whether the questions can be supported by valid or sufficient supporting rationales, some of the questions in these benchmarks may be inherently noninferable. To address this issue, we introduce a new benchmark, PROPHET, which comprises inferable forecasting questions paired with relevant news for retrieval. To ensure the inferability of the benchmark, we propose Causal Intervened Likelihood (CIL), a statistical measure that assesses inferability through causal inference. In constructing this benchmark, we first collected recent trend forecasting questions, and then filtered the data using CIL resulting in an inferable benchmark for event prediction. Through extensive experiments, we first demonstrate the validity of CIL  and in-depth investigations into event prediction with the aid of CIL. Subsequently, we evaluate several representative prediction systems on PROPHET, drawing valuable insights for future directions. The dataset is available on https://github.com/TZWwww/PROPHET.
PROPHET: An Inferable Future Forecasting Benchmark with Causal Intervened Likelihood Estimation
Zhengwei Tao12thanks: 🖂 Corresponding author.   Zhi Jin12🖂  Bincheng Li3  Xiaoying Bai4🖂  Haiyan Zhao12   Chengfeng Dou12   Xiancai Chen12   Jia Li12   Linyu Li12   Chongyang Tao5 1Key Laboratory of High Confidence Software Technologies (PKU), MOE, China 2School of Computer Science, Peking University  3Guangzhou University,  4Advanced Institute of Big Data  5SKLSDE Lab, Beihang University {tttzw,xiancaich}@stu.pku.edu.cn,  {zhijin,zhhy.sei,chengfengdou,lijiaa}@pku.edu.cn baixy@aibd.ac.cn,  chongyang@buaa.edu.cn
1 Introduction
@@
The quest to predict future events has long been a central pursuit in the field of artificial intelligence (AI). The ability to foresee outcomes and trends holds the promise of revolutionizing numerous sectors covering finance Li et al. (2024), climate science Wang and Karimi (2024), and social policy Rotaru et al. (2022). Recent years have witnessed a surge in interest and progress, particularly with the advent of large language model (LLM)-based systems. These systems, leveraging the power of deep learning and vast amounts of data, have demonstrated an unprecedented capacity for forecasting, capturing the imagination and focus of the research community Halawi et al. (2024); Hsieh et al. (2024); Pratt et al. (2024).
@@
Figure 1: The upper Figure demonstrates the task of future forecasting. The lower half shows both inferable and non-inferable scenarios.
@@
To evaluate the abilities of these LLM-based future forecasting systems, pilot works construct several benchmarks based on real-world forecasting questions Halawi et al. (2024); Guan et al. (2024); Karger et al. (2024). These benchmarks have successfully framed future forecasting as a retrieval-augmented generation (RAG)-and-reasoning task. Within this framework, systems should first search the Web or databases for news articles related to the prediction question in the benchmarks to gain knowledge base, then reason based on the retrieved knowledge base. Nevertheless, in order to truly evaluate the abilities of the LLM-based future forecasting, the prediction questions in the benchmarks need to be inferable, meaning that the supporting knowledge base must contain sufficient information to substantiate the answers. In traditional RAG tasks, the answer can definitely be found within the knowledge base. However, future forecasting tasks do not inherently satisfy this characteristic compared to traditional RAG benchmarks such as HotpotQA Yang et al. (2018) and 2WikiMultiHopQA Ho et al. (2020). That is, future forecasting needs to be inferred by rationales, i.e. facts and reasoning clues, but the knowledge base may only provide partially supportive rationales for the prediction questions Zhao et al. (2024). Collecting real-world prediction questions as the benchmark without nuanced validation, the knowledge base may not be able to provide sufficient supportive facts which makes some of the prediction questions non-inferable Birur et al. (2024).
@@
To overcome this challenge and advance the field, we introduce an inferable future forecasting benchmark, PROPHET, designed to provide a more accurate evaluation. To ensure reproducibility, PROPHET is an RAG task where each prediction question pairs with relevant downloaded news articles for retrieval. We are next motivated to select prediction questions that are inferable, based on their related articles. The most challenging part is to estimate the inferability of each question since we cannot observe the completed real-world event evolution process. Even if we can, it is difficult to determine as well, due to the lack of expert knowledge of a wide spectrum of domains. A key innovation in our approach is the introduction of Causal Intervened Likelihood (CIL), a statistical measure that assesses the inferability of prediction questions through causal inference. CIL is calculated via principles of causal inference where we measure the supporting degree of each article for the answer to the question. We regard each article as an event and compute the effect of intervening in the event from happening to not happening. CIL provides a robust estimate of whether a question can be answered. We then filter the prediction questions using CILto ensure the inferability of the benchmark, providing a fair and accurate evaluation of the systems’ forecasting ability. Assisted by CIL, PROPHET performs as a more well-formulated RAG-and-reasoning task with hidden rationale Zhao et al. (2024).
@@
    We are the first to introduce CIL for inferability estimation of the future forecasting questions and provide a feasible method for calculating this metric.
@@
    Assisted by CIL, we establish an automatic pipeline to construct the future forecasting benchmark PROPHET where the prediction questions are insufficiently inferable based on their related articles.
@@
2 Related Work
2.1 Future Forecasting and Benchmarks
Previous research on future forecasting benchmarks has evolved in different paradigms, each addressing different aspects of the task. Early benchmarks, such as MCNC Granroth-Wilding (2016), SCT Mostafazadeh et al. (2017), and CoScript Yuan et al. (2023), focused on script learning and common sense reasoning in synthetic scenarios. Although these data sets facilitated structured reasoning, they lacked real-world applicability and grounding in factual news. Time series datasets such as GDELT Leetaru and Schrodt (2013) and ICEWS Schrodt et al. (2012) introduced real-world event tracking but did not formalize prediction as a retrieval-augmented reasoning task or ensure answerability. Later works, such as ECARE Du et al. (2022) and EV2 Tao et al. (2024), advanced event reasoning but remained confined to settings without real-world grounding.
With the rise of LLMs, recent benchmarks such as Halawi et al. (2024), OpenEP Guan et al. (2024), and ForecastBench Karger et al. (2024) shifted the focus to real-world questions and news-based search. However, these datasets suffer from two critical limitations: (1) they lack explicit validation of inferability, allowing questions with insufficient supporting evidence to persist, and (2) they prioritize dynamic data sources over reproducibility, risking inconsistent evaluations due to evolving news archives. PROPHET  addresses these gaps by filtering via the introduced Causal Intervened Likelihood estimation. We show the benchmark comparison in Table 4.
2.2 RAG and Benchmarks
Foundational QA Datasets for RAG: Traditional QA datasets, including MMLU Hendrycks et al. (2021), StrategyQA Geva et al. (2021), ASQA Stelmakh et al. (2022), Multi-HopQA Lin et al. (2020), and 2WikiMultiHopQA Lin et al. (2020), are adapted to evaluate RAG systems. These datasets, grounded in knowledge bases like Wikipedia, form the basis for RAG evaluation.
Domain-Agnostic: RAGBench Friel et al. (2024) is a multi-domain benchmark across biomedical, legal, customer support, and finance domains. CRAG Wang et al. (2024a) provides a factual QA benchmark across five domains, simulating web and knowledge graph search.
Domain-Specific: Domain-specific benchmarks include LegalBench-RAG Wang et al. (2024b), WeQA Meyur et al. (2024), PubHealth Zhang et al. (2023), and MTRAG Tang and Yang (2024). These benchmarks address niche applications and improve evaluation precision in domains.
Capability-Oriented: RGB Liu et al. (2024) evaluates four RAG capabilities: noise robustness, negative rejection, information integration, and counterfactual robustness. TRIAD Zong et al. (2024) assesses retrieval quality, fidelity, and task-specific utility through a three-dimensional framework.
3 Preliminaries
3.1 Future Forecasting
@@
Future forecasting stands for predicting whether a certain event will happen in the future based on the events that occurred. We now formalize the task as a binary question-answering task. Given a prediction question ${\mathcal{Q}}$ which can be “Will Tim Walz win the VP debate against J.D. Vance?” or “Will Bitcoin rise to \$100,000 by December 2024?”. There would be background information ${\mathcal{B}}$ that describes the context of ${\mathcal{Q}}$ and resolution criteria ${\mathcal{R}}$ explaining how the question can be regarded as answered. A large set of documents ${\mathbb{X}}$ serves as a knowledge base to retrieve. The forecasting system must answer the question as:
${\mathcal{Y}}=\mathrm{Reason}({\mathcal{Q}},{\mathcal{B}},{\mathcal{R}},\mathrm{Retrieve}({\mathcal{Q}},{\mathbb{X}})),$ (1)
where ${\mathcal{Y}}\in[0,1]$ is the predicted probability of how likely the event in ${\mathcal{Q}}$ would occur. A ground truth answer $\hat{{\mathcal{Y}}}\in\{0,1\}$ paired with a resolved date ${\mathcal{D}}$ represents whether the event in ${\mathcal{Q}}$ finally occurs and the date the question resolves. As the same in previous works Halawi et al. (2024); Karger et al. (2024), we use Brier Score Brier (1950) as the metric for evaluation:
Brier $\texttt{Brier Score}=\frac{1}{N}\sum_{n}^{N}({\mathcal{Y}}_{n}-\hat{{\mathcal{Y}}}_{n})^{2},$ (2)
$N$ is the number of the questions in the dataset.
@@
We formalize future forecasting as an RAG task. As an RAG, it features distinctly compared with traditional dataset such as HotpotQA Yang et al. (2018) and 2WikiMultiHopQA Ho et al. (2020). The knowledge base ${\mathbb{X}}$ stores the rationales and clues for answering ${\mathcal{Q}}$ Zhao et al. (2024). Future forecasting mainly detects two core entangled abilities of the systems: retrieval and reasoning.
@@
Current future forecasting benchmarks are constructed by harvesting real-world prediction questions and paired with news articles before the resolved date ${\mathcal{D}}$ Halawi et al. (2024); Guan et al. (2024); Karger et al. (2024) without nuanced validation of the inferability of the questions. It is possible that there is a lack of sufficient supportive information in ${\mathbb{X}}$ for the question. Methods need to be established to ensure that the prediction questions in the benchmarks are sufficiently inferable.
3.2 Causal Inference
@@
Causal inference is a vital statistical method to determine causal relationships between variables Pearl (2010). In real-world scenarios, a mere correlation between two variables may be due to chance or hidden factors. Causal inference aims to establish direct causality. For example, the increase in ice cream sales and drowning incidents is not a causal link, although both are affected by hot weather. Causal inference uses concepts such as structural causal models, interventions, and counterfactual inferences. These are applied in medicine, economics, and social sciences.
@@
Structural causal model (SCM) It is a framework designed to represent and analyze causal relationships between variables using a combination of causal graphs and structural equations. At its core, SCM relies on a directed graph where nodes represent variables ${\mathcal{X}}$, and edges denote direct causal influences, forming a network that captures dependencies and pathways of causation. Each variable in the model is determined by its direct causes (parent nodes). SCM enables the identification of causal effects, and exploration of intervention questions (e.g., "What would happen if we intervened on X?"). This has been widely applied in fields like epidemiology, economics, and machine learning to disentangle complex causal mechanisms and validate hypotheses Stolfo et al. (2023).
@@
Interventional distribution An SCM allows the study of interventions. An atomic intervention $\text{{do}}({\mathcal{X}}_{i}=x)$ fixes ${\mathcal{X}}_{i}$ with a fixed value $x$. For example, in a medical trial, the dose of a new drug is set at a specific value for a group. In the view of structural causal model, interventions can be understand as changing of the original structure and variable distributions. After $\text{{do}}({\mathcal{X}}_{i}=x)$, the resulting distribution is $P(\cdot|\text{{do}}({\mathcal{X}}_{i}=x))\doteq P_{{m}}(\cdot|{\mathcal{X}}_{i}=x)$, which shows how other variables respond.
4 PROPHET Benchmark
@@
In this section, we introduce PROPHET  which is an future forecasting benchmark with inferability estimation and selection. We first describe the data collection process in Section 4.1. Then we introduce the Causal Intervened Likelihood (CIL) metric in Section 4.2. We finally describe the benchmark construction in Section 4.3.
4.1 Data Collection
@@
Our objective is to gather a dataset that encompasses recent and prominent prediction questions. To achieve this, we have sourced questions from two well-known platforms: Metaculas [^1] https://www.metaculus.com and Manifold [^2] https://manifold.markets. The choice of these source websites, Metaculas and Manifold, is well justified for constructing the benchmark. The domains covered by the questions on these platforms are highly diverse, ranging from scientific breakthroughs to social and economic trends. This diversity ensures that the benchmark is representative of a wide spectrum of forecasting tasks. Moreover, the questions are trending and among the most attention-attracting ones on these platforms. This indicates that they are not only relevant in the current context but also likely to be of interest to the broader forecasting community. As such, the data collected from these sources provides a robust foundation for evaluating and developing practical forecasting models.
To avoid model leakage, we carefully selected questions. From Metaculas, we chose questions resolved in August 2024 along with meta-information. Since there were few pre-August 2024 questions on Metaculas, we added questions resolved before August from Manifold, ensuring both the latest trends and a historical perspective. We filtered out meaningless questions, such as personal inquiries or those with little community interest, to focus on realistic forecasting scenarios.
@@
After collecting questions, we collected relevant news articles. Using GPT4o-mini [^3] https://openai.com, we generated three types of news search queries per question: entities in the question, resolving steps, and similar historical events using prompts in the Appendix A.7 (a-c). Then we searched on the MediaCloud open-source platform [^4] https://www.mediacloud.org with these queries. MediaCloud’s vast news repository helped us gather comprehensive information. However, many retrieved articles were irrelevant. To address this, we used GPT4o-mini again to filter the articles, retaining 100 relevant ones per question by prompt in the Appendix A.7 (d). That reduces noise and mimics real-world prediction analysis.
4.2 Causal Intervened Likelihood
To measure the sufficiency of the supportive rationales of each question and construct an inferable benchmark, we introduce a statistic estimation named Causal Intervened Likelihood (CIL) via causal inference. CIL estimates the supportivity of each news article to the question. We use Bernoulli variables to model the occurrence of events. Specifically, let $Y\in\{0,1\}$ indicate whether the event asked by the question happens or not, and let $X_{i}\in\{0,1\}$ indicate whether the situation described in the $i$-th news happens or not. Each variable ${\mathcal{X}}_{i}$ is associated with a date ${\mathcal{T}}_{i}$ since each news also has the occurrence date. We use the notation ${\mathcal{T}}_{i}\prec{\mathcal{T}}_{j}$ to represent that the occurrence of the $i^{th}$ news is before that of the $j^{th}$. Note that the date of ${\mathcal{Y}}$ is later than any date of ${\mathcal{X}}$.
Intuitively, if the $i^{th}$ news article’s occurrence ($X_{i}=1$) constitutes a necessary condition for $Y=\hat{Y}$ (ground-truth answer), then the intervention $\mathrm{do}(X_{i}=0)$ would significantly increase the probability of $Y\neq\hat{Y}$. With this intuition, we define the CIL of the $i^{th}$ news article as:
$\texttt{CIL}_{i}=$ $P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=1))$ (3)
$-P({\mathcal{Y}}=\hat{{\mathcal{Y}}}|do({\mathcal{X}}_{i}=0)),$
where $do$ is the intervention operation in causal inference standing for ${\mathcal{X}}$ is intervened to happen or not as stated in Section 3.2.
To compute this estimation, we model all ${\mathcal{X}}_{i}$ and ${\mathcal{Y}}$ as a structural causal model (SCM). For this SCM, we treat all ${\mathcal{X}}_{i}$ and ${\mathcal{Y}}$ as nodes and causal relationships between them as edges. However, it is extremely hard to extract causal edges in our case due to incomplete knowledge base and intensive dependency on experts. It is difficult to calculate CIL via methods relying on the complete SCM.
To fill this gap, we introduce three assumptions. We illustrate these assumptions in Figure 2. Firstly, the causal relations between the news should be aligned with temporality. This assumption is consistent with common sense and eliminates circle paths in the SCM. Notice that ${\mathcal{Y}}$ is the variable in this SCM with the latest date.
@@ 4.2 Causal Intervened Likelihood
Figure 2: Illustration of assumptions. Nodes represent news variables that are in chronological order corresponding to their ${\mathcal{T}}$.

### v2

@@ Abstract.
Predicting future events based on news on the Web stands as one of the ultimate aspirations of artificial intelligence. Recent advances in large language model (LLM)-based systems have shown remarkable potential in forecasting future events, thereby garnering significant interest in the research community. Currently, several benchmarks have been established to evaluate the forecasting capabilities by formalizing the event prediction as a retrieval-augmented generation (RAG)-and-reasoning task. In these benchmarks, each prediction question is answered with relevant retrieved news articles downloaded from the Web. However, because there is no consideration of whether the questions can be supported by valid or sufficient supporting rationales, some of the questions in these benchmarks may be inherently noninferable. To address this issue, we introduce a new benchmark, PROPHET, which comprises inferable forecasting questions paired with relevant news for retrieval. To ensure the inferability of the benchmark, we propose Causal Intervened Likelihood (CIL), a statistical measure that assesses inferability through causal inference. In constructing this benchmark, we first collected recent trend forecasting questions, and then filtered the data using CIL resulting in an inferable benchmark for future forecasting. Through extensive experiments, we first demonstrate the validity of CIL  and in-depth investigations into future forecasting with the aid of CIL. Subsequently, we evaluate several representative prediction methods on PROPHET. The overall results draws valuable insights for task of future directions. Data is public on https://github.com/TZWwww/PROPHET.
copyright: acmlicensedjournalyear: 2018doi: XXXXXXX.XXXXXXXconference: Preprint; January, 2026;
1. Introduction
@@
The quest to forecast future events based on information on the Web has long been a central pursuit in the field of artificial intelligence (AI). The ability to foresee outcomes and trends holds the promise of revolutionizing numerous sectors covering finance (Li et al., 2024), climate science (Wang and Karimi, 2024), and social policy (Rotaru et al., 2022). Recent years have witnessed a surge in interest and progress, particularly with the advent of large language model (LLM)-based systems. These systems, leveraging the power of deep learning and vast amounts of data, have demonstrated an unprecedented capacity for forecasting, capturing the imagination and focus of the research community (Halawi et al., 2024; Hsieh et al., 2024; Pratt et al., 2024).
@@
Figure 1. The upper Figure demonstrates the task of future forecasting. The lower half shows both inferable and non-inferable scenarios.
@@
To evaluate the abilities of these LLM-based future forecasting systems, pilot works construct several benchmarks based on real-world forecasting questions (Halawi et al., 2024; Guan et al., 2024; Karger et al., 2024; Zeng et al., 2025). These benchmarks have successfully framed future forecasting as a retrieval-augmented generation (RAG)-and-reasoning task. Within this framework, systems should first search the Web or databases for news articles related to the prediction question in the benchmarks to gain knowledge base, then reason based on the retrieved knowledge base. Nevertheless, in order to truly evaluate the abilities of the LLM-based future forecasting, the prediction questions in the benchmarks need to be inferable, meaning that the supporting knowledge base must contain sufficient information to substantiate the answers. In traditional RAG tasks, the answer can definitely be found within the knowledge base. However, future forecasting tasks do not inherently satisfy this characteristic compared to traditional knowledge-intensive benchmarks such as HotpotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020). That is, future forecasting needs to be inferred by rationales, i.e. facts and reasoning clues, but the knowledge base may only provide partially supportive rationales for the prediction questions (Zhao et al., 2024). Collecting real-world prediction questions as the benchmark without nuanced validation, the knowledge base may not be able to provide sufficient supportive facts which makes some of the prediction questions non-inferable (Birur et al., 2024).
@@
To overcome this challenge, we introduce an inferable future forecasting benchmark, PROPHET, designed to provide a more accurate evaluation. To ensure reproducibility, PROPHET is an RAG task where each real-world prediction question pairs with relevant downloaded news articles for retrieval from the Web. We are next motivated to select prediction questions that are inferable, based on their related articles. The most challenging part is to estimate the inferability of each question since we cannot observe the completed real-world event evolution process. Even if we can, it is difficult to determine as well, due to the lack of expert knowledge of a wide spectrum of domains. A key innovation in our approach is the introduction of Causal Intervened Likelihood (CIL), a statistical measure that assesses the inferability of prediction questions through causal inference. CIL is calculated via principles of causal inference where we measure the supporting degree of each article for the answer to the question. We regard each article as an event and compute the effect of intervening in the event from happening to not happening. CIL provides a robust estimate of whether a question can be answered. We then filter the prediction questions using CIL to ensure the inferability of the benchmark, providing a fair and accurate evaluation of the systems’ forecasting ability. Assisted by CIL, PROPHET performs as a well-formulated RAG-and-reasoning task with hidden rationale (Zhao et al., 2024).
@@
    We are the first to introduce CIL for inferability estimation of the future forecasting. We provide a feasible method for calculating this metric, which we also testify to its validity..
@@
    Assisted by CIL, we establish an automatic pipeline to construct the future forecasting benchmark PROPHET where the real-world prediction questions are insufficiently inferable based on their related news articles.
@@
2. Preliminaries
2.1. Future Forecasting
@@
Future forecasting stands for predicting whether a certain event will happen in the future based on news information from the Web. We now formalize the task as a binary question-answering task. Given a prediction question $Q$ which can be “Will Tim Walz win the VP debate against J.D. Vance?” or “Will Bitcoin rise to \$100,000 by December 2024?”. There would be background information $B$ that describes the context of $Q$. A large set of news articles $\mathbb{X}$} serves as a knowledge base to retrieve. The forecasting system must answer the question as formalized:
(1) $Y=\mathrm{Reason}(Q,B,{\mathbb{X}}),$
[…]
(2) $\texttt{Brier Score}=\frac{1}{N}\sum_{n}^{N}(Y_{n}-\hat{Y}_{n})^{2},$
NN is the number of questions in the dataset. In this work, we report the Brier Score expanded by 100 times and keep 2 decimal places.
@@
We formalize future forecasting as an RAG task. As an RAG, it features distinctly compared with traditional dataset such as HotpotQA (Yang et al., 2018) and 2WikiMultiHopQA (Ho et al., 2020). The knowledge base $\mathbb{X}$} stores the rationales and clues for answering $Q$ (Zhao et al., 2024). Future forecasting mainly detects two core entangled abilities of the systems: retrieval and reasoning.
@@
Current future forecasting benchmarks are constructed by harvesting real-world prediction questions and paired with news articles before the resolved date $D$ (Halawi et al., 2024; Guan et al., 2024; Karger et al., 2024) without nuanced validation of the inferability of the questions. It is possible that there is a lack of sufficient supportive information in $\mathbb{X}$} for the question. Methods need to be established to ensure that the prediction questions in the benchmarks are sufficiently inferable.
2.2. Causal Inference
@@
Causal inference is a vital statistical method to determine causal relationships between variables (Pearl, 2010). In real-world scenarios, a mere correlation between two variables may be due to chance or hidden factors. Causal inference aims to establish direct causality. For example, the increase in ice cream sales and drowning incidents is not a causal link, although both are affected by hot weather. Causal inference uses concepts such as structural causal models, interventions, and counterfactual inferences. These are applied in medicine, economics, and social sciences.
@@
Structural causal model (SCM) It is a framework designed to represent and analyze causal relationships between variables using a combination of causal graphs and structural equations. At its core, SCM relies on a directed graph where nodes represent variables $X$, and edges denote direct causal influences, forming a network that captures dependencies and pathways of causation. Each variable in the model is determined by its direct causes (parent nodes). SCM enables the identification of causal effects, and the exploration of intervention questions (e.g., ”What would happen if we intervened on X?”). This has been widely applied in fields like epidemiology, economics, and machine learning to disentangle complex causal mechanisms and validate hypotheses (Stolfo et al., 2023).
@@
Interventional distribution An SCM allows the study of interventions. An atomic intervention $\text{{do}}(X_{i}=x)$ fixes $X_{i}$ with a fixed value $x$. For example, in a medical trial, the dose of a new drug is set at a specific value for a group. In the view of the structural causal model, interventions can be understand as changing of the original structure and variable distributions. After $\text{{do}}(X_{i}=x)$, the resulting distribution is $P(\cdot|\text{{do}}(X_{i}=x))\doteq P_{{m}}(\cdot|X_{i}=x)$, which shows how other variables respond.
A critical distinction in causal inference lies between interventional and observational probabilities. The observational probability, denoted as the conditional probability $P(Y|X=x)$, represents the likelihood of event $Y$ given that we have passively observed variable $X$ to be $x$. This probability captures mere statistical association or correlation. In contrast, the interventional probability, denoted as $P(Y|\text{do}(X=x))$, represents the likelihood of $Y$ if we were to actively intervene and set the value of $X$ to $x$. This distribution describes the causal effect of $X$ on $Y$ by simulating a controlled experiment and removing spurious correlations. While interventional probabilities are essential for predicting the effects of actions, they cannot be computed directly from passively collected data. Therefore, a primary challenge in causal inference is to determine whether and how these interventional probabilities can be calculated from available observational probabilities(Pearl, 2010).
3. PROPHET Benchmark
@@
In this section, we introduce PROPHET  which is a future forecasting benchmark with inferability estimation and selection. We first describe the data collection process in Section 3.1. Then we introduce the Causal Intervened Likelihood (CIL) metric in Section 3.2. We describe the benchmark construction in Section 3.4 and  3.3. We report the statistics of PROPHET in Section 3.5.
3.1. Data Collection
@@
Our objective is to gather a dataset that encompasses recent and prominent prediction questions. To achieve this, we have sourced questions from well-known platforms: Polymarket [^1] https://polymarket.com/. The domains covered by the questions on these platforms are highly diverse, ranging from scientific breakthroughs to social and economic trends. This diversity ensures that the benchmark is representative of a wide spectrum of forecasting tasks. Moreover, the questions are trending and among the most attention-attracting ones on the platform. This indicates that they are not only relevant in the current context but also likely to be of interest to the broader forecasting community. As such, the data collected from these sources provides a robust foundation for evaluating and developing practical forecasting models.
@@
To avoid model leakage, we carefully selected questions. We aim to use the latest forecast questions. We include all questions on Polymarket resolved between 2025,1,1 and 2025,1,31. We filter out meaningless questions, such as personal inquiries or those with little community interest, to focus on realistic forecasting scenarios. After collecting questions, we collected relevant news articles as much as possible. Using LLM, we generated three types of news search queries per question: entities in the question, resolving steps, and similar historical events using prompts in the Appendix Prompts (a-c). Then we searched on the MediaCloud open-source platform [^2] https://www.mediacloud.org with these queries. MediaCloud’s vast news repository helped us gather comprehensive news URLs. After gathering the news URLs, we use Newspaper3k [^3] https://newspaper.readthedocs.io/ API to download the news articles. We discard news that is later than the resolved date of the question.
However, many downloaded articles were irrelevant. There are also news articles tagged with wrong publish date, which could incur data leakage. To address these, we use LLM to rate the relevant score of each article to the question. We set the highest score, indicating this article directly answering of the question, while the lowest score showing total irrelevance. We remove both articles with the highest and lowest scores. The relevance rating prompt is in the Appendix Prompts (d).
@@
Figure 2. Illustration of assumptions. Nodes are variables that are in chronological order with time $T$.
# Question # News Avg Article Word Avg CIL Avg Pos CIL Avg Neg CIL Earlies News Date Latest News Date
612 559.51 1249.88 -0.013 0.2041 -0.2503 2023,1,1 2025,1,1
Table 1. Statistics of L1 part of PROPHET.
3.2. Causal Intervened Likelihood
To measure the sufficiency of supportive rationales for each question and to construct an inferable benchmark, we introduce a statistical estimation named Causal Intervened Likelihood (CIL), derived from causal inference. CIL estimates how strongly each news article supports the answer to a given question.
We use Bernoulli variables to model the occurrence of events. Let $Y\in\{0,1\}$ indicate whether the event asked by the question occurs, and let $X_{i}\in\{0,1\}$ indicate whether the event described in the $i$-th news article occurs. Each variable $X_{i}$ is associated with a happening date $T_{i}$. We use the notation $T_{i}\prec T_{j}$ to represent that the occurrence of the $i$-th news is before that of the $j$-th. Note that the date of $Y$ is later than any date of $X_{i}$.
Intuitively, if the occurrence of the $i$-th news article ($X_{i}=1$) is a necessary condition for the answer $Y=\hat{Y}$, then the intervention $\mathrm{do}(X_{i}=0)$ would significantly increase the probability of $Y\neq\hat{Y}$. With this intuition, we define the CIL of the $i$-th news article as:
(3) $\texttt{CIL}_{i}=P(Y=\hat{Y}|\mathrm{do}(X_{i}=1))-P(Y=\hat{Y}|\mathrm{do}(X_{i}=0)),$
where $\mathrm{do}(\cdot)$ is the intervention operation from causal inference, representing that $X_{i}$ is forced to happen or not.
The CIL estimation is composed by interventional probabilities. As similar to the core idea in causal inference, we derive these interventional probabilities to observational probabilities. We model all $X_{i}$ and $Y$ as a Structural Causal Model (SCM), where variables are nodes and causal relationships are directed edges. However, determining the complete SCM is extremely challenging due to incomplete knowledge and the need for intensive expert analysis. Therefore, calculating CIL via methods that rely on a complete SCM is impractical. Thus we introduce two key assumptions.